OpenStudy (anonymous):

Simplify (2-3i) -(6+4i) a) -24-10i b) 4+7i c) -4-7i d) -4+7i

OpenStudy (cwrw238):

2 - 6 - 3i - 4i simplify like terms

OpenStudy (cwrw238):

can u finish this?

OpenStudy (cwrw238):

2 - 6 = ?

OpenStudy (anonymous):

-4?

OpenStudy (cwrw238):

right and -3i - 4i ?

OpenStudy (anonymous):

not sure? either +7 or -7

OpenStudy (cwrw238):

well - don't forget the 'i' -3i - 4i = -7i

OpenStudy (anonymous):

So it's c?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

ok thanks

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

Do you know how to do this one? Give the value of the discriminantof the equation 4x^-7x+3=0 a) 1 b)-128 c)0 d)128

OpenStudy (cwrw238):

you might be wondering why -(6 + 4i) = -6 - 4i think of it being -1(6 + 4i) = -1*6 + -1* 4i = -6 - 4i

OpenStudy (cwrw238):

i expect you mean 4x^2-7x+3=0 for ax^2 + bx + c = 0 discriminant = b^2 - 4ac for your equation b = -7 , a = 4 and c = 3 plug these into the discriminant formula to get its value

OpenStudy (cwrw238):

= (-7)^2 - 4*4*3

OpenStudy (anonymous):

So would it be a?

OpenStudy (cwrw238):

= 1 yes a

OpenStudy (anonymous):

Factor 8x^2+14x+3 Show Work

OpenStudy (cwrw238):

sorry - gotta go